Consumption 445 +0.75y Investment 250 Calculate the value of autonomous spending Calculate the value of multiplier Calculate the Equilibrium level of income Main determination of the level of investment
1
Expert's answer
2021-10-25T08:44:03-0400

1.Autonomous spending

A= total consumption that is independent of income

C=445+0.75 x 0

A= 445

2.Calculate the value of multiplier

Multiplier= 1/(1-MPC)

From the consumption function, MPC=0.75

Multiplier = 1/(1-0.75)

Multiplier = 1/0.25

Multiplier= 4


3.Calculate the Equilibrium level of income

at equilibrium, Y= C+I

Y= 445+0.75Y+250

Y-0.75Y=445+250

0.25Y= 695

Y= 695/0.25

Y= 2780


4.Main determination of the level of investment

Availability of finance

Interest rates

Economic growth

Confidence

Technology